OpenStudy (anonymous):

Simplify the following expression. x - 4y + 5z - 3x - y + 2z A) 2x - 5y + 7z B) -2x - 5y - 7z C) -2x - 5y + 7z D) -2x - 5y + 3z

OpenStudy (anonymous):

@Notamathgenius

OpenStudy (anonymous):

A's out

OpenStudy (anonymous):

really? i thought it was C! lol but okay

OpenStudy (notamathgenius):

B is wrong -_-

OpenStudy (anonymous):

leaves only computing \(2z+5z\)

OpenStudy (anonymous):

it's C

OpenStudy (notamathgenius):

Wait B is wright , my bad :P

OpenStudy (anonymous):

no B is wrong

OpenStudy (anonymous):

So B? everybody agree on B?

OpenStudy (anonymous):

no B is wrong

OpenStudy (anonymous):

okay....

OpenStudy (anonymous):

\[2z+5z=?\]

OpenStudy (notamathgenius):

You want to simplify, So that means putting z with z y with y x with x and any others that are equal together.

OpenStudy (anonymous):

then what?

OpenStudy (notamathgenius):

After thats its simplified.

OpenStudy (anonymous):

oh okay lol so C?

OpenStudy (notamathgenius):

yes :D

OpenStudy (anonymous):

okay c: Solve: 4/5 y + 2 =6 5 7 9 11

OpenStudy (anonymous):

dont answer yet, i think i got this.

OpenStudy (notamathgenius):

haha ok :D

OpenStudy (anonymous):

its A lol sorry i had to find my glasses .-. im blindddd

OpenStudy (notamathgenius):

Me too :D

OpenStudy (anonymous):

cool n_n

OpenStudy (anonymous):

near or far? im far c:

OpenStudy (notamathgenius):

No i mean like a wall will appear in front of me without warning -_-

OpenStudy (anonymous):

lmao xD i did that once with a tree... i was driving on a dirt road .-. i love how tough ford broncos are n_n

OpenStudy (notamathgenius):

hahaha XDD

OpenStudy (anonymous):

it was sugar sands fault though... it jerked me and i ran OVER the tree .-. it was pretty big too lol

OpenStudy (notamathgenius):

Im only 15 i ride dirtbikes though that is fun when you hit thinks haha that sucks XD

OpenStudy (anonymous):

im 15 too :p but i live in the middle of nowhere, nobody notices :p

OpenStudy (notamathgenius):

Way to go lol

OpenStudy (anonymous):

Choose the graph below that correctly represents the equation 3x - 4y = 12. (pictures coming up)

OpenStudy (anonymous):

A)

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

C)

OpenStudy (anonymous):

D)

OpenStudy (notamathgenius):

Grab a pair of Coords and plug in for X and Y, but that will take to long the Answer is B ;P

OpenStudy (anonymous):

lol thank you C:

OpenStudy (notamathgenius):

No problem :D

OpenStudy (anonymous):

A cell phone plan offers a monthly rate of $25 for unlimited minutes plus $.05 per text message. Which algebraic expression below represents the total cost for this cell phone plan in terms of "x?" 25.05x 25x + 0.05 25 + 0.05x 24.95x

OpenStudy (anonymous):

C?

OpenStudy (notamathgenius):

Not A B or D what must it be? ;D

OpenStudy (anonymous):

i got it right then c;

OpenStudy (notamathgenius):

haha yes :P

OpenStudy (anonymous):

Let A = {circle, square} and let B = {3, 4, 5}. Find A X B. A) {circle, square, 3, 4, 5} B) { } C) {3, 4, 5} D) {(circle, 3), (circle, 4), (circle, 5), (square, 3), (square, 4), (square, 5)}

OpenStudy (notamathgenius):

Ok this one may look hard but its not, I can tell you B and A and C are wrong.

OpenStudy (anonymous):

so D? lol

OpenStudy (notamathgenius):

D.5 :P yes D

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 2 > -3 and 3x <= 12 A) A number line with an open circle on -1, shading to the left, and a closed circle on 4, shading to the right. B) A number line with an open circle on -1, a closed circle on 4, and shading in between. C) A number line with a closed circle on -1, shading to the left, and an open circle on 4, shading to the right. D) A number line with a closed circle on -1, an open circle on 4, and shading in between.

OpenStudy (notamathgenius):

Hmmm, well I hate these a lot.

OpenStudy (anonymous):

meee tooo

OpenStudy (notamathgenius):

Can you guess, I am not wanting to give you a wrong answer. D:

OpenStudy (anonymous):

okay C:

OpenStudy (notamathgenius):

:D

OpenStudy (anonymous):

2 + (a + b) = (2 + a) + b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

i forgot .-.

OpenStudy (notamathgenius):

Its B or C.

OpenStudy (anonymous):

Evaluate: -3 - (-6) -3 -18 -9 3

OpenStudy (notamathgenius):